OpenStudy (cggurumanjunath):

\int { mg\cos { \theta } d\theta } =\int { mR\ddot { \theta } } d\theta

OpenStudy (cggurumanjunath):

\[\int\limits { mg\cos { \theta } d\theta } =\int\limits { mR\ddot { \theta } } d\theta \]

OpenStudy (cggurumanjunath):

integrate wrt theta ,theta dot and theta double dot. i will give medal .can aby body help.

OpenStudy (anonymous):

What math is this? This looks really hard.

OpenStudy (cggurumanjunath):

integration @asdfasdf54514

OpenStudy (lukecrayonz):

@satellite73

OpenStudy (ash2326):

\[\int mg \cos \theta d\theta\] mg are constant with respect to theta, bring them out \[mg\int \cos \theta d\theta\] can you integrate it now? @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

my question is can u integtate lhs and rhs wrt theta ? @ash2326

OpenStudy (anonymous):

is it the original question ?

OpenStudy (cggurumanjunath):

s

OpenStudy (ash2326):

yes, you can do that

OpenStudy (cggurumanjunath):

can u show me the steps or draw which ever is convenient to u @ash2326

OpenStudy (unklerhaukus):

the m's will cancel

OpenStudy (ash2326):

\[\int\limits { mg\cos { \theta } d\theta } =\int\limits { mR\ddot { \theta } } d\theta\] \[mg\int \cos \theta d\theta=mR\int \ddot \theta d\theta\] We know that \[\frac{d}{d\theta} \dot \theta=\ddot \theta\] so \[mg\int \cos \theta d\theta=mR\dot \theta +C\] where C is constant, I think you can integrate the left side, can't you?

OpenStudy (anonymous):

the dots are wrt to t or theta ?!

OpenStudy (anonymous):

usually dots mean derivative with respect to time

OpenStudy (cggurumanjunath):

ok @ash2326 u have integrated wrt theta right ?

OpenStudy (cggurumanjunath):

can u integrate wrt theta dot.

OpenStudy (ash2326):

yes, cause we have dtheta in the equation. No info about time

OpenStudy (cggurumanjunath):

no

OpenStudy (cggurumanjunath):

no information is given wrt time.

OpenStudy (anonymous):

@CGGURUMANJUNATH are the dots mean derivative with respect to theta or with respect to time ?

OpenStudy (cggurumanjunath):

theta @Coolsector

OpenStudy (anonymous):

ok.

OpenStudy (cggurumanjunath):

can u show me the steps plz @Coolsector

OpenStudy (anonymous):

what steps ? if you say that the dots are with respect to theta then see the solution ash gave you

OpenStudy (cggurumanjunath):

ok what ash has done wrt theta is fine ,can u integrate wrt theta dot.

OpenStudy (cggurumanjunath):

@Coolsector

OpenStudy (cggurumanjunath):

theta dot =\dot { \theta }

OpenStudy (cggurumanjunath):

\[\dot { \theta } \]

OpenStudy (cggurumanjunath):

i meant.

OpenStudy (cggurumanjunath):

my question is can u integtate lhs wrt theta and rhs wrt theta dot ;is it possible to do that ?

OpenStudy (anonymous):

you mean in case your original equation: mgcos(theta) = mR* (theta double dot ) ?

OpenStudy (anonymous):

if so you cant integrate like this one side wrt to one varible and the other wrt to another

OpenStudy (cggurumanjunath):

thank you very much @Coolsector @ash2326 and @UnkleRhaukus

OpenStudy (anonymous):

anyway. i still cant get it. if dot means derivative with respect to theta then theta dot should be 1.

OpenStudy (anonymous):

it is like taking the derivative of x wrt to x. and the second derivative would be 0.

OpenStudy (cggurumanjunath):

it is like taking the derivative of x wrt to x.YES IT IS RIGHT, then theta dot should be 1. HOW @coolsector.

OpenStudy (cggurumanjunath):

@Coolsector @ash2326 @UnkleRhaukus chech out the attachment.

OpenStudy (cggurumanjunath):

sorry check not chech this was the actual problem .

OpenStudy (unklerhaukus):

well the derivative shold be with respect to time as @Coolsector sugessted.

OpenStudy (cggurumanjunath):

ok.